# SongGen: A Single Stage Auto-regressive Transformer for Text-to-Song Generation

### v1

@@ 4.3 Ablation Studies
Effect of Lyrics Module Design. We further investigate the impact of different lyric integration methods, including the choice of tokenizer (VoiceBPE vs. T5), the use of a lyrics encoder, and the integration approach (pre-pending vs. cross-attention). Figure 4 shows the results after Step 1 training for each variant. Our design (VoiceBPE, w/ lyrics encoder, cross-attention) achieves the best results across all metrics, validating the effectiveness. Unlike most TTS works, which prepend transcripts before audio tokens, we find that the cross-attention approach is more effective and stable. This may be because cross-attention allows the decoder to focus solely on generating the audio modality. Additionally, phoneme-like tokenizer (VoiceBPE) is more suitable than the word-level tokenizer (T5) for song lyric tokenization. Under this mechanism, the lyrics encoder can capture the relationships between lyric tokens, learning pronunciation patterns from different token combinations, and thus alleviating the burden of modality alignment on the decoder.
@@ 5 Conclusion
In conclusion, while SongGen presents exciting possibilities for the music industry and creative expression, its development should be accompanied by careful consideration of its ethical and societal implications
Appendix A Limitations and Future Work
Appendix B Implementation and Training Details
In SongGen, the lyrics encoder is a 6-layer transformer with a hidden size of 1024. The transformer decoder, consisting of 24 layers with 1024 hidden size, includes both causal self-attention and cross-attention blocks in each layer. In “Mixed Pro” Mode, the vocal loss weight $\lambda$ is set to 0.1. The model is trained for approximately 400K steps using 16 Nvidia A100 (80GB) GPUs, with a batch size of 16 per GPU. For optimization, we employ the AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a weight decay of $10^{-4}$. During training step 1, the learning rate is set to $10^{-4}$, while for the subsequent fine-tuning steps, the learning rate is reduced to $5\times 10^{-5}$. A cosine learning rate schedule is applied for all traning steps. To facilitate reproducibility, we will make our training configurations publicly available.
Appendix C Details in Evaluations
@@
For objective metrics, all samples are normalized at-14dB LUFS for fairness. We use the following five metrics:
  • Frechet Audio Distance (FAD) (Kilgour et al., 2019): Evaluates the fidelity of generated songs by calculating the distribution distance between features of the target and generated audio, extracted from the VGGish (Hershey et al., 2017) model.
  • Kullback-Leibler Divergence (KL): Measures the similarity between the generated and target audio with the label calculated by the audio tagging model. A lower KL suggests that the generated music shares similar concepts with the reference.
  • CLAP Score: Evaluates the alignment between generated audio and the given text prompt using the official CLAP model (Wu* et al., 2023).
  • Phoneme Error Rate (PER): Assesses the adherence of the generated audio to the provided lyrics by transcribing the audio using Distill Whisper(Gandhi et al., 2023) and computing the phoneme error rate against the reference lyrics. However, PER is not an ideal measure of vocal quality, as current ASR models struggle with sung vocals.
  • Speaker Embedding Cosine Similarity (SECS): Assesses the similarity of speaker identity using the Resemblyzer [^3] Implemented based on: https://github.com/resemble-ai/Resemblyzer speaker encoder to compute the SECS between reference 3-second vocal clips and generated audio.
@@
For the subjective evaluations, we randomly select 36 audio samples generated by our models, and each sample is evaluated by 20 listeners. We conduct the commonly used MOS (Mean Opinion Score) tests across five aspects. The rating scale ranges from 1 to 5, with higher scores indicating better performance. For the Overall Quality (OVL.) evaluation, we instruct the raters to focus on musicality and naturalness, while ignoring style differences. For the Relevance to Text Description (REL.) evaluation, we ask the raters to score based on the proportion of key points from the text description that are reflected in the generated song. For the Vocal Quality (VQ.) evaluation, we emphasize the importance of clarity, lyric accuracy, and the naturalness and coherence of the vocals in the ratings. For Harmony (HAM.), we ask the raters to pay particular attention to the temporal correspondence between the accompaniment and the vocals. For Speaker Similarity, we ask the raters to focus on the similarity of the speaker’s identity (timbre) to the reference, ignoring differences in content. A small subset of the samples used in the test is available on our anonymous project page https://liuzh-19.github.io/SongGen/.
Appendix D The Impact of Different Audio Codecs.
@@
We compare the performance of three different codecs: XCodec, Encodec (24kHz) (Défossez et al., 2022), and DAC (44.1kHz) (Kumar et al., 2024). X-Codec, employed in our SongGen, considers both acoustic and semantic information. In contrast, Encodec and DAC have been widely used in previous audio generation systems, particularly in speech (Lyth & King, 2024) and pure music domains (Copet et al., 2024), but both primarily focus on acoustic reconstruction. Table 5 shows the results after training Step 1. X-Codec greatly surpasses both Encodec and DAC on all metrics. Additionally, the loss curves in Figure 6 demonstrate that X-Codec exhibits more stable training and faster convergence. This indicates that the integration of semantic information into the audio token is highly effective and crucial for the song generation task.

### v2

@@ 4.3 Ablation Studies
Effect of Lyrics Module Design.
We further investigate the impact of different lyric integration methods, including the choice of tokenizer (VoiceBPE vs. T5), the use of a lyrics encoder, and the integration approach (pre-pending vs. cross-attention). Figure 5 shows the results after Step 1 training for each variant. Our design (VoiceBPE, w/ lyrics encoder, cross-attention) achieves the best results across all metrics, validating the effectiveness. Unlike most TTS works, which prepend transcripts before audio tokens, we find that the cross-attention approach is more effective and stable. This may be because cross-attention allows the decoder to focus solely on generating the audio modality. Additionally, phoneme-like tokenizer (VoiceBPE) is more suitable than the word-level tokenizer (T5) for song lyric tokenization. Under this mechanism, the lyrics encoder can capture the relationships between lyric tokens, learning pronunciation patterns from different token combinations, and thus alleviating the burden of modality alignment on the decoder.
5 Limitations and Future Work
@@ 6 Conclusion
Acknowledgments
This work was supported by National Key R&D Program of China 2022ZD0161600, Shanghai Artificial lntelligence Laboratory, the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK. Dahua Lin is a PI of CPII under the InnoHK.
@@ Acknowledgments
In conclusion, while SongGen presents exciting possibilities for the music industry and creative expression, its development should be accompanied by careful consideration of its ethical and societal implications.
Appendix A Implementation and Training Details
In SongGen, the lyrics encoder is a 6-layer transformer with a hidden size of 1024. The transformer decoder, consisting of 24 layers with 1024 hidden size, includes both causal self-attention and cross-attention blocks in each layer. In “Mixed Pro” Mode, the vocal loss weight $\lambda$ is set to 0.1. The model is trained for approximately 400K steps using 16 Nvidia A100 (80GB) GPUs, with a batch size of 16 per GPU. For optimization, we employ the AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a weight decay of $10^{-4}$. During training step 1, the learning rate is set to $10^{-4}$, while for the subsequent fine-tuning steps, the learning rate is reduced to $5\times 10^{-5}$. A cosine learning rate schedule is applied for all traning steps. The resource allocation and training strategy for the multi-stage baseline are consistent with those used for SongGen mixed training Step 1. Since voice-free support is not directly related to the core comparison between the single-stage and multi-stage designs, we omit this part of the multi-stage model. Specifically, the two models in the multi-stage pipeline are trained separately for approximately 200K steps each. We observe that the loss begins to plateau around 60K steps for both models. To facilitate reproducibility, we make our training configurations publicly available.
Appendix B Details in Evaluations
@@
For objective metrics, all samples are normalized at-14dB LUFS for fairness. Frechet Audio Distance (FAD) (Kilgour et al., 2019) evaluates the fidelity of generated songs by calculating the distribution distance between features of the target and generated audio, extracted from the VGGish (Hershey et al., 2017) model. Kullback-Leibler Divergence (KL) measures the similarity between the generated and target audio with the label calculated by the audio tagging model. A lower KL suggests that the generated music shares similar concepts with the reference. FAD and KL are calculated using audioldm_eval  [^4] https://github.com/haoheliu/audioldm_eval. CLAP Score evaluates the alignment between generated audio and the given text prompt using the official CLAP model (Wu* et al., 2023), implemented via the stable-audio-metrics  [^5] https://github.com/Stability-AI/stable-audio-metrics. However, we observe that the CLAP Score for MusicCaps test set is unexpectedly low. We supplement our evaluation with the CLaMP3 Score (Wu et al., 2025), a more recent model that provides a more robust measure of text–song alignment. Phoneme Error Rate (PER) assesses the adherence of the generated audio to the provided lyrics by transcribing the audio using Distill Whisper(Gandhi et al., 2023) and computing the phoneme error rate against the reference lyrics. However, PER is not an ideal measure of vocal quality, as current ASR models struggle with sung vocals. Speaker Embedding Cosine Similarity (SECS) assesses the similarity of speaker identity using the Resemblyzer [^6] https://github.com/resemble-ai/Resemblyzer speaker encoder to compute the SECS between reference 3-second vocal clips and generated audio. We further introduce recently proposed content-based aesthetics metrics (Tjandra et al., 2025), covering Content Enjoyment (CE), Content Usefulness (CU), Production Complexity (PC), and Production Quality (PQ).
@@
For the subjective evaluations, we randomly select 36 audio samples generated by our models, and each sample is evaluated by 20 listeners. We conduct the commonly used MOS (Mean Opinion Score) tests across five aspects. The rating scale ranges from 1 to 5, with higher scores indicating better performance. For the Overall Quality (OVL.) evaluation, we instruct the raters to focus on musicality and naturalness, while ignoring style differences. For the Relevance to Text Description (REL.) evaluation, we ask the raters to score based on the proportion of key points from the text description that are reflected in the generated song. For the Vocal Quality (VQ.) evaluation, we emphasize the importance of clarity, lyric accuracy, and the naturalness and coherence of the vocals in the ratings. For Harmony (HAM.), we ask the raters to pay particular attention to the temporal correspondence between the accompaniment and the vocals. For Speaker Similarity, we ask the raters to focus on the similarity of the speaker’s identity (timbre) to the reference, ignoring differences in content. A small subset of the samples used in the test is available on our project page https://liuzh-19.github.io/SongGen/.
Appendix C The Impact of Different Audio Codecs.
@@
We compare the performance of three different codecs: XCodec, Encodec (24kHz) (Défossez et al., 2022), and DAC (44.1kHz) (Kumar et al., 2024). Table 6 shows the results after training Step 1. X-Codec surpasses both Encodec and DAC on all metrics. Additionally, the loss curves in Figure 6 demonstrate that X-Codec exhibits more stable training and faster convergence. Although Encodec and DAC have been widely adopted in prior audio generation systems across domains such as speech (Lyth & King, 2024) and instrumental music (Copet et al., 2024), song generation presents a substantially higher level of semantic complexity. While Encodec and DAC indeed yield better perceptual quality for audio reconstruction, we observed that in song generation, both codecs resulted in higher rates of invalid outputs, such as failure to follow lyrics, or producing noise and silence. In contrast, X-Codec consistently demonstrated more stable training, faster convergence, and higher success rates in generating coherent vocals.
We attribute this performance to several factors. First, we adopt the publicly released xcodec_hubert_general_audio checkpoint, trained on a large-scale (200k-hour) private dataset with a distribution similar to AudioSet (Gemmeke et al., 2017). We speculate that its exposure to large amounts of music data during pretraining contributes to its superior performance in our task. Second, as emphasized in the X-Codec paper (Ye et al., 2024), the incorporation of not only acoustic but also semantic features from self-supervised learning representations might also contribute to the performance. Despite X-Codec operating at a relatively low sampling rate of 16 kHz, we selected it as the most suitable option available at the time. To date, high-fidelity, song-specific neural codecs tailored for generative modeling remain an open challenge in the research community.